OpenStudy (anonymous):

Differential Equations Help! Undetermined coefficients y''+3y=-9

OpenStudy (abb0t):

start by finding the complimentary solution first, then your particular.

OpenStudy (zzr0ck3r):

what is the characteristic equation?

OpenStudy (anonymous):

is my general solution \[y=C1Cos(2\sqrt{3}x)+C2Sin(2\sqrt{3}x)\]

OpenStudy (abb0t):

\(r^2+3 = 0\) \(\Rightarrow\) \(r= \pm \sqrt{3}i\) so,

OpenStudy (anonymous):

oh damn I noob'd it and used -9 :/

OpenStudy (abb0t):

wait, how did you get \(2\sqrt{3}\)???

OpenStudy (abb0t):

anyways, i THINK they say to use a "guess" for your constant, but I am more familiar with the annihilator method, which is somewhat the same thing.

OpenStudy (anonymous):

instead of r^2+3=0 I did r^2+3=-9

OpenStudy (anonymous):

yeah my bad thanks though cause I probably would've never found it

OpenStudy (abb0t):

oh. ok. so you're ok from here or did you still need more help?

OpenStudy (anonymous):

im gonna try to work it if you'd like to give me a few min ill put up what I get or where I get stuck

OpenStudy (anonymous):

does A=-3?

OpenStudy (anonymous):

@abb0t if you get a chance to check my answer it would be greatly appreciated!

OpenStudy (abb0t):

right.

OpenStudy (anonymous):

beast! thanks